# Self-Training Elicits Concise Reasoning in Large Language Models

### v2

@@ 1 Introduction
We posit that current models often generate more tokens than necessary to accomplish the task, incurring extraneous inference costs. Typical reasoning chains sampled from the original model distribution, exemplified in the blue box of Figure 1, includes many tokens that do not contribute to the solution, specifically verbose explanations and repetitive phrasing. Similar observations have also been made in prior work Renze and Guven (2024); Zhang et al. (2024); Chiang and Lee (2024a).
@@
In this paper, we propose a simple yet effective fine-tuning method to elicit efficient, concise reasoning in models for a given target task. We note that existing zero-shot prompting methods fails to reliably elicit concise reasoning, exhibiting near negligible impact on task-specialized models (Section 2.2). On the other hand, we can leverage best-of-N (BoN) sampling and few-shot conditioning (FS) to reliably generate concise training data (Section 3). We can then apply standard fine-tuning to distill the length-reducing benefits of BoN sampling and few-shot prompting back into the model itself while avoiding their associated overheads at inference time.
Across GSM8K and MATH datasets and a wide variety of model families, our unified few-shot conditioned best-of-N sampling (FS-BoN) method achieves a substantial reduction in output length of 30% on average, a 2.4x improvement over previous fine-tuning baselines De Sabbata et al. (2024), while preserving overall accuracy. Notably, our analysis shows that trained models adaptively adjust output length based on question complexity, preserving detail for difficult problems while simplifying responses to easier ones. We confirm that these results remain consistent across various model scales. These results strongly suggest that fine-tuning with carefully curated, self-generated data effectively can unlock latent concise reasoning abilities within LLMs, leading to significantly more cost-effective inference for complex tasks.
@@ 5.1 Main results
FS-BoN elicits the largest length reduction among our self-training methods, while maintaining relative accuracy, on average. Notably, for math-specialized models, FS-GPT4o-BoN achieves the greatest reduction among all methods, except those fine-tuned on external data which greatly sacrifice the accuracy (Table 8 and Table 9).
@@ 6 Discussion
In this paper, we focused on standard fine-tuning based on self-generated samples. Our analysis shows that LLMs already possess the ability to correctly reason in a relatively concise manner. Therefore, we posit that lightweight fine-tuning is sufficient to achieve significant reasoning length reduction, echoing the Superficial Alignment Hypothesis in the field of LLM alignment Zhou et al. (2023), Indeed, our post-hoc analysis in Figure 7 shows that standard fine-tuning can reduce the output length of models in proportion to that of training samples.
@@ 7 Conclusion
We tackle redundancy in CoT reasoning, hypothesizing LLMs possess a latent capacity for concise reasoning, evidenced by shorter correct reasoning paths. We introduce fine-tuning methods, leveraging self-generated data from BoN sampling and few-shot conditioning, to elicit this capacity. Our FS-BoN method significantly reduces reasoning length by 30% reduction while maintaining accuracy. This implies fine-tuning with curated self-generated data can reliably unlock latent concise reasoning, enabling more efficient inference.
@@
Although we leveraged few-shot prompting to bootstrap length reduction during data generation, we did not explore advanced few-shot prompting methods as a middle ground between zero-shot and fine-tuning. While direct few-shot prompting with our prompts achieved comparable length reduction to fine-tuning on single-path few-shot generations, it incurred a slight accuracy loss. Critically, fine-tuning enabled us to incorporate BoN sampling and sample augmentation, leading to significantly greater length reduction without sacrificing accuracy. However, future work on integrating advanced exemplar selection Fu et al. (2022) and many-shot prompting Agarwal et al. (2024) could further enhance our fine-tuning approach.
@@ A.2 Concise Reasoning
Many studies have shown that LLMs tend to overthink during CoT, often including unnecessary words in their reasoning (Zhang et al., 2024; Chiang and Lee, 2024b; Wang et al., 2024), which can negatively impact efficiency. However, simply reducing the number of tokens often leads to significantly lower performance  (Merrill and Sabharwal, 2023; Jin et al., 2024). To address this, some works have explored prompting techniques to elicit concise reasoning (Renze and Guven, 2024; Nayab et al., 2024; Han et al., 2024), but these methods are highly model-dependent. Another line of work has focused on fine-tuning methods to balance efficiency and reasoning quality (De Sabbata et al., 2024), but these approaches are insufficient in reducing token usage. In this work, we investigate how to efficiently generate concise reasoning samples and use them for fine-tuning, improving model efficiency without sacrificing accuracy.
Several concurrent works also address redundancies in LLM outputs. Some approaches Chen et al. (2024); Luo et al. (2025) focus on how o1-like models overthink and propose methods for pruning extraneous reasoning steps. Another approach, TokenSkip Xia et al. (2025), first identifies and discards low-importance tokens from self-generated CoT outputs while preserving crucial ones, then fine-tunes the model on these compressed examples to enable efficient inference. In contrast, our method elicits concise reasoning more simply yet effectively by training on a self-generated dataset of shorter rationales, using best-of-N sampling and few-shot conditioning.
@@ B.2 Generation of Few-Shot Exemplars From GPT-4o (FS-GPT4o)
To obtain few-shot exemplars of concise reasoning from GPT-4o, we use our zero-shot prompt Hand Crafted 3 (Appendix F.1) to generate short reasoning paths. For GSM8K, we randomly sample eight questions from the training set. For MATH, we randomly select two questions from the largest category, Algebra and one from each of the other categories (Geometry, Intermediate Algebra, Prealgebra, Number Theory, Counting & Probability, Precalculus), all from the training set. We manually filter low-quality outputs to ensure the quality of few-shot examples. However, we found that GPT-4o reliably followed the zero-shot prompt in the majority of cases. Generated examples of GPT-4o are in Tables 15 and 16.
@@ Appendix E Sample Study
Tables 10-13 present additional examples of generated reasoning paths from both the initial model and the model fine-tuned using FS-GPT4o-BoN. We primarily showcase results for the general model Llama-3.2-3b and the math-fine-tuned model Deepseek-7b, evaluated on the GSM8k and MATH500 datasets. One can observe a consistent reduction in token length in our fine-tuned models, achieved through decreased verbosity. This demonstrates the effectiveness of our method in improving inference efficiency.
@@ F.2 Few-Shot Prompts
We provide 8-shot examples for each question, following the approach of Wei et al. (2022b), without any system instructions given to the language model. Each example is formatted in a chat-based template that includes both user and assistant roles. The order of examples for each question is randomly shuffled. We present one representative example for each prompt type, while the full set of examples is available in our public repository. Table 14-18 display one of eight examples used in FS-Human, FS-GPT4o, and FS-Self.

### v3

@@ 1 Introduction
We posit that current models often generate more tokens than necessary to accomplish the task, incurring extraneous inference costs. Typical reasoning chains sampled from the original model distribution, exemplified in the blue box of Figure 1, includes many tokens that do not contribute to the solution, specifically verbose explanations and repetitive phrasing. Similar observations have also been made in prior work Renze and Guven (2024); Zhang et al. (2024); Chiang and Lee (2024).
@@
In this paper, we propose a simple yet effective fine-tuning method to elicit efficient, concise reasoning in models for a given target task. We note that existing zero-shot prompting methods fail to reliably elicit concise reasoning, exhibiting near negligible impact on task-specialized models (Section 2.2). Instead, we can leverage best-of-N (BoN) sampling and few-shot conditioning (FS) to reliably generate concise training data (Section 3). We can then apply standard fine-tuning to distill the length-reducing benefits of BoN sampling and few-shot prompting back into the model itself while avoiding inference-time overheads.
Across GSM8K and MATH datasets and a wide variety of model families, our unified few-shot conditioned best-of-N sampling (FS-BoN) method achieves a substantial reduction in output length of 30% on average, a 2.4x improvement over previous fine-tuning baselines De Sabbata et al. (2024), while preserving overall accuracy. Our analysis shows that trained models adaptively adjust output length based on question complexity, preserving detail for difficult problems while simplifying responses to easier ones. We confirm that these results remain consistent across various model scales and diverse reasoning domains such as business, chemistry, and physics. This shows that fine-tuning with carefully curated, self-generated data can effectively unlock latent concise reasoning abilities within LLMs, leading to significantly more cost-effective inference for complex tasks.
@@ 5.1 Main results
FS-BoN elicits the largest length reduction among our self-training methods, while maintaining relative accuracy, on average. Notably, for math-specialized models, FS-GPT4o-BoN achieves the greatest reduction among all methods, except those fine-tuned on external data which greatly sacrifice the accuracy (Table 13 and Table 14). Moreover, robustness testing with different few-shot exemplar sets confirms consistent performance, maintaining accuracy while effectively reducing reasoning length even with varied exemplars (Section D.9).
@@
Cross-task evaluation demonstrates broader applicability.
We conducted cross-task evaluation by testing models fine-tuned on one dataset against another. Our method shows strong generalization, achieving 10-12% length reductions with minimal accuracy drops when applied to out-of-domain mathematical problems (detailed analysis in Section D.7).
@@
Broader domain evaluation demonstrates consistent effectiveness.
We evaluate our method on diverse reasoning domains (business, chemistry, physics) from MMLU-Pro (Wang et al., 2024a) that benefit from chain-of-thought reasoning (Sprague et al., 2024) in Appendix D.8. Our method demonstrates consistent effectiveness across all three domains, achieving both accuracy improvements (16.51% average) and substantial length reduction (26.82% average). Notably, our method provides over 3x better length reduction efficiency compared to naive BoN while maintaining competitive accuracy gains.
@@ 5.2 Analysis
Practical efficiency gains
We also conducted wall-clock time and memory usage measurements after fine-tuning models with our method. The results show that latency is reduced by 15.38% to 52.94%, while memory usage decreases by 2.50% to 6.26% across two different models (detailed analysis provided in Section D.6). These improvements directly translate to lower computational costs and faster inference in real-world deployments.
@@ 6 Discussion
In this paper, we focused on standard fine-tuning based on self-generated samples. Our analysis shows that LLMs already possess the ability to correctly reason in a relatively concise manner. Therefore, we posit that lightweight fine-tuning is sufficient to achieve significant reasoning length reduction, echoing the Superficial Alignment Hypothesis in the field of LLM alignment Zhou et al. (2023). Indeed, our post-hoc analysis in Figure 7 shows that standard fine-tuning can reduce the output length of models in proportion to that of training samples.
@@ 7 Conclusion
We tackle redundancy in CoT reasoning by hypothesizing that LLMs possess a latent capacity for concise reasoning, as evidenced by shorter correct reasoning paths within their output distributions. We introduce fine-tuning methods that leverage self-generated data from best-of-N sampling and few-shot conditioning to elicit this capacity. Our FS-BoN method significantly reduces reasoning length by 30% while maintaining accuracy. This implies fine-tuning with curated self-generated data can reliably unlock latent concise reasoning, enabling more efficient inference.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.2019-0-00075, Artificial Intelligence Graduate School Program (KAIST) and No. RS-2024-00457882, AI Research Hub Project) and Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT(MSIT, Korea) & Gwangju Metropolitan City.
@@ Acknowledgements
Although we leveraged few-shot prompting to bootstrap length reduction during data generation, we did not explore advanced few-shot prompting methods as a middle ground between zero-shot and fine-tuning. While direct few-shot prompting with our prompts achieved comparable length reduction to fine-tuning on single-path few-shot generations, it incurred a slight accuracy loss. Critically, fine-tuning enabled us to incorporate BoN sampling and sample augmentation, leading to significantly greater length reduction without sacrificing accuracy. Moreover, our robustness analysis across different exemplar sets confirms consistent performance. However, future work on integrating advanced exemplar selection Fu et al. (2022) and many-shot prompting Agarwal et al. (2024) could further enhance our fine-tuning approach.
@@ A.2 Concise Reasoning
Many studies have shown that LLMs tend to overthink during CoT, often including unnecessary words in their reasoning (Zhang et al., 2024; Chiang and Lee, 2024; Wang et al., 2024b), which can negatively impact efficiency. However, simply reducing the number of tokens often leads to significantly lower performance  (Merrill and Sabharwal, 2023; Jin et al., 2024). To address this, some works have explored prompting techniques to elicit concise reasoning (Renze and Guven, 2024; Nayab et al., 2024; Han et al., 2024), but these methods are highly model-dependent. Another line of work has focused on fine-tuning methods to balance efficiency and reasoning quality (De Sabbata et al., 2024), but these approaches are insufficient in reducing token usage. In this work, we investigate how to efficiently generate concise reasoning samples and use them for fine-tuning, improving model efficiency without sacrificing accuracy.
Several concurrent works also address redundancies in LLM outputs. Some approaches Chen et al. (2024); Luo et al. (2025); Team et al. (2025) focus on how o1-like models overthink and propose methods for pruning extraneous reasoning steps. Another approach, TokenSkip Xia et al. (2025), first identifies and discards low-importance tokens from self-generated CoT outputs, then fine-tunes the model on these compressed examples to enable efficient inference. In contrast, our method elicits concise reasoning more simply yet effectively by training on a self-generated dataset of shorter rationales, using best-of-N sampling and few-shot conditioning.
@@ B.2 Generation of Few-Shot Exemplars From GPT-4o (FS-GPT4o)
To obtain few-shot exemplars of concise reasoning from GPT-4o, we use our zero-shot prompt Hand Crafted 3 (Appendix F.1) to generate short reasoning paths. For GSM8K, we randomly sample eight questions from the training set. For MATH, we randomly select two questions from the largest category, Algebra and one from each of the other categories (Geometry, Intermediate Algebra, Prealgebra, Number Theory, Counting & Probability, Precalculus), all from the training set. We manually filter low-quality outputs to ensure the quality of few-shot examples. However, we found that GPT-4o reliably followed the zero-shot prompt in the majority of cases. Generated examples of GPT-4o are in Tables 20 and 21.
@@ D.4 Full Result Tables on Main Methods
Dataset (# samples) Model Base (sec) Ours (sec) ↓%
GSM8K (1319) Ll-3B 16 10 37.50
GSM8K (1319) QM-1.5B 17 8 52.94
MATH (500) Ll-3B 17 13 23.53
MATH (500) QM-1.5B 13 11 15.38
Table 7: Wall-clock time comparison between baseline zero-shot prompting and our method (FS-GPT4o-BoN) using vLLM on a single H100 GPU for two models: the general-purpose Llama-3.2-3B and the task-specific Qwen2.5-Math-1.5B.
@@ D.5 Impact of Augmentation on Training Data Length
D.6 Efficiency Analysis: Wall-Clock Time and Memory Usage
We conduct additional efficiency measurements to quantify the practical benefits of our method (FS-GPT4o-BoN) beyond token count reduction.
Wall-Clock Time Savings
We measured wall-clock generation times using vLLM, a highly optimized implementation, on a single H100 GPU. Our results confirm that models trained with our method (FS-GPT4o-BoN) achieve significant wall-clock time reductions ranging from 15.38% to 52.94% compared to the original models using baseline zero-shot prompting. The results are presented in Table 7.
Dataset (# samples) Model Base (MB) Ours (MB) ↓%
GSM8K (1319) Ll-3B 17873.76 16756.52 6.25
GSM8K (1319) QM-1.5B 6250.06 6044.87 3.28
MATH (500) Ll-3B 34739.72 33871.16 2.50
MATH (500) QM-1.5B 13019.13 12342.20 5.20
Table 8: Peak memory usage comparison between baseline zero-shot prompting and our method (FS-GPT4o-BoN) using HuggingFace with batch size 128 on a single H100 GPU for two models: the general-purpose Llama-3.2-3B and the task-specific Qwen2.5-Math-1.5B.
Memory Usage Reduction
We also measured memory usage to assess efficiency gains. Since vLLM automatically utilizes all available memory, we compared memory usage under fixed batch-size decoding using HuggingFace, a popular LLM framework. Running inference with a batch size of 128 on a single H100 GPU, we found that peak memory usage (measured via torch.cuda.max_memory_allocated) was reduced by 2.50% to 6.25% after fine-tuning the model with our method. These results are shown in Table 8.
D.7 Cross-Task Generalization Analysis
We conducted cross-task evaluation to assess how well our method generalizes beyond task-specific fine-tuning. We fine-tuned models on one dataset and evaluated them on another across all five main models (Llama-3.2-3B, Gemma-2-2B, Qwen2.5-3B, Qwen2.5-Math-1.5B, DeepSeekMath-7B).
Train Test Rel Acc (%) Rel Len (%)
GSM8K In-Domain (GSM8K) 97.00 64.25
Out-of-Domain (MATH) 99.21 88.34
MATH In-Domain (MATH) 102.56 76.30
Out-of-Domain (GSM8K) 98.53 90.36
Table 9: Cross-task evaluation results using FS-GPT4o-BoN method. Models are fine-tuned on one dataset and evaluated on another, averaged across five main models. Relative metrics are computed against the baseline performance on each respective evaluation dataset.
Our results show that reasoning length is reduced even under out-of-domain (OOD) evaluation, with only minor drops in accuracy as shown in Table 9. For example, when training on GSM8K and testing on MATH, we observed a 0.79% relative drop in accuracy alongside an 11.66% reduction in output length. Conversely, when training on MATH and testing on GSM8K, we found a 1.47% relative drop in accuracy with a 9.64% reduction in output length. These results suggest that our method generalizes reasonably well across different types of math problems.
D.8 Evaluation of Broader Reasoning Domains
To demonstrate that our method is applicable beyond mathematical domains, we conducted additional experiments on diverse subject areas that benefit from chain-of-thought reasoning. We evaluated our method on the business, chemistry, and physics categories from MMLU-Pro (Wang et al., 2024a), which have been shown to benefit significantly from CoT reasoning compared to direct answering (Sprague et al., 2024).
We trained and evaluated three general-purpose models (Llama-3.2-3B, Qwen2.5-3B, Gemma-2-2B) on these MMLU-Pro categories using a random 7:3 train-test split. We compared our FS-GPT4o-BoN method against naive BoN and baseline zero-shot prompting, following the same methodology described in Section 3.
Domain FS-GPT4o-BoN Naive BoN
Rel Acc (%) Rel Len (%) Rel Acc (%) Rel Len (%)
Business 102.78 67.38 113.03 92.48
Chemistry 116.33 78.25 119.48 92.99
Physics 130.43 73.90 124.79 91.02
Average 116.51 73.18 119.10 92.16
Table 10: Evaluation results on broader reasoning domains from MMLU-Pro. Results are averaged across three general-purpose models (Llama-3.2-3B, Qwen2.5-3B, Gemma-2-2B) using a random 7:3 train-test split. Relative metrics are computed against baseline zero-shot prompting.
Our results show that reasoning length is reduced while accuracy is increased compared to baseline zero-shot prompting across all domains, as shown in Table 10. Compared to naive BoN, our FS-GPT4o-BoN method achieves more than 3× relative length reduction, though with a slightly smaller increase in accuracy.
D.9 Robustness to Few-Shot Exemplars
Dataset Examples Ll-3B QM-1.5B
Acc Length Acc Length
GSM8K Baseline 77.26 220.00 82.79 305.91
Original 77.79 119.89 80.74 124.56
New 1 77.03 114.23 80.36 125.04
New 2 78.09 128.93 82.11 146.60
MATH Baseline 44.40 511.75 69.80 551.36
Original 43.60 316.12 67.40 378.64
New 1 42.80 363.39 69.40 429.96
New 2 45.80 373.91 68.40 408.63
Table 11: Robustness analysis of FS-GPT4o-BoN with different few-shot exemplar sets. Results show accuracy (%) and reasoning length (tokens) across two models and datasets. "Original" refers to the exemplars used in main experiments, while "New 1" and "New 2" are randomly selected alternative sets.
To validate the robustness of our few-shot conditioning approach, we conducted additional experiments with different sets. We re-evaluated our FS-GPT4o-BoN method using two new sets of few-shot examples, randomly selected from the training set following the same generation procedure described in Section B.2. We tested on two representative models: Llama-3.2-3B (general-purpose) and Qwen2.5-Math-1.5B (math-specialized), across both GSM8K and MATH datasets.
Our results show that FS-GPT4o-BoN maintains accuracy while effectively reducing reasoning length even with different few-shot exemplars, further demonstrating the robustness of our approach as shown in Table 11.
@@ Appendix E Sample Study
Tables 15-18 present additional examples of generated reasoning paths from both the initial model and the model fine-tuned using FS-GPT4o-BoN. We primarily showcase results for the general model Llama-3.2-3b and the math-fine-tuned model Deepseek-7b, evaluated on the GSM8k and MATH500 datasets. One can observe a consistent reduction in token length in our fine-tuned models, achieved through decreased verbosity. This demonstrates the effectiveness of our method in improving inference efficiency.
@@ F.2 Few-Shot Prompts
We provide 8-shot examples for each question, following the approach of Wei et al. (2022b), without any system instructions given to the language model. Each example is formatted in a chat-based template that includes both user and assistant roles. The order of examples for each question is randomly shuffled. We present one representative example for each prompt type, while the full set of examples is available in our public repository. Table 19-23 display one of eight examples used in FS-Human, FS-GPT4o, and FS-Self.